OpenStudy (anonymous):

A new Firestone tire is guaranteed to last for 40,000 miles. The actual mean life of the tires is 47,000 miles with a standard deviation of 4,000 miles

OpenStudy (anonymous):

What are you doing here? Testing a hypothesis? Finding some probability?

OpenStudy (anonymous):

what is the probability that a tire will not last 40,000 miles

OpenStudy (anonymous):

\[P(X<40,000)=P\left(\frac{X-47,000}{4,000}<\frac{40,000-47,000}{4,000}\right)=P(Z<-1.75)\] Consult a \(z\) table, like the one here: http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf